set up an equation and answer the following

A man is 4 times as older as her daughter . After 16 years , he will be twice as old as her daughter age . find daughter 's age

Answer by mananth(16946)

daughter =x
father =4x
After 16 years
daughter = x+16
father = 4x+16
4x+16 = 2(x+16)
4x+16=2x+32
4x-2x=32-16
2x=16
x=8
daughter = 8 years
father = 32 years